# EVF-SAM: Early Vision-Language Fusion for Text-Prompted Segment Anything Model

### v3

@@ Abstract
Segment Anything Model (SAM) has attracted widespread attention for its superior interactive segmentation capabilities with visual prompts while lacking further exploration of text prompts. In this paper, we empirically investigate what text prompt encoders (e.g., CLIP or LLM) are good for adapting SAM for referring expression segmentation and introduce the Early Vision-language Fusion-based SAM (EVF-SAM). EVF-SAM is a simple yet effective referring segmentation method which exploits multimodal prompts (i.e., image and text) and comprises a pre-trained vision-language model to generate referring prompts and a SAM model for segmentation. Surprisingly, we observe that: (1) multimodal prompts and (2) vision-language models with early fusion (e.g., BEIT-3) are beneficial for prompting SAM for accurate referring segmentation. Our experiments show that the proposed EVF-SAM based on BEIT-3 can obtain state-of-the-art performance on RefCOCO/+/g for referring expression segmentation and demonstrate the superiority of prompting SAM with early vision-language fusion. In addition, the proposed EVF-SAM with 1.32B parameters achieves remarkably higher performance while reducing nearly 82% of parameters compared to previous SAM methods based on large multimodal models. The code, demo and models are available at https://github.com/hustvl/EVF-SAM.
Index Terms:
Referring image segmentation, vision-language models, multimodal models, segment anything.
@@ 1 Introduction
Segment Anything Model (SAM) [1] brings interactive segmentation paradigm to public view. Well-trained on the SA-1B dataset, SAM achieves stunning performance and quickly becomes popular as a vision foundation model for object localization and beyond. Various SAM variants [2, 3, 4, 5] have been explored, achieving better efficiency or higher precision. Despite SAM’s surprising abilities like point-prompted and box-prompted segmentation, it is a pity that the text-prompted segmentation ability remains conceptual. We retrospect such task to Referring Expression Segmentation (RES). RES focuses on the solution that one predicts the segmentation mask according to the text description given by users, which enjoys several explorations by some traditional end-to-end models [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19], and is broadened by some Large Multimodal Models (LMM) [20, 21, 22, 23, 24, 25, 26, 27].
@@
The key challenge lies in empowering SAM with language understanding ability for segmentation according to text prompts, e.g., referring expression segmentation. Fig. 2 summarizes previous works which explore the text-prompted abilities of SAM: (a) SAM with grounded detector: A two-stage framework where a grounded detector generates a bounding box to prompt SAM, e.g., Grounded-SAM [28]. However, those methods suffer from a sub-optimal architecture, where segmentation heavily relies on the accuracy of the detector, and it is difficult to optimize due to its non-end-to-end nature. (b) SAM with text encoder: A off-the-shelf text encoder, e.g., CLIP [29], is used to encode the text prompt, providing text embeddings for SAM. Whereas the semantic gap exists between the text embeddings and SAM which is pre-trained with geometric prompts, i.e., points or boxes, thus the segmentation performance is inferior. (c) SAM with LLM: A Large Language Model (LLM) (or Large Multimodal Model) is employed and fine-tuned to get the desired embeddings about object information. The embeddings will be used to predict segmentation masks based on image features. However, these LLM-based models are often computationally expensive, requiring massive memory and computation budgets, and the training is challenging. Additionally, complex conversation templates need to be manually designed to instruct the LLM for referring segmentation. Can we leverage a more efficient but effective method to empower SAM with text-prompted ability in an end-to-end manner?
@@
Motivated by the above observations, we extend SAM for language understanding and text-prompt capabilities by incorporating a Multimodal Encoder with Early Vision-Language Fusion (EVF) and present EVF-SAM in this paper. The proposed EVF-SAM aims to be a simple framework to prompt SAM with texts and illustrate how to prompt SAM to follow referring expressions effectively. EVF-SAM is built on the off-the-shelf foundation models and comprises a Multimodal Encoder, an early-fused vision-language model, e.g., BEIT-3 [30], and a simple projector to generate prompt embeddings for SAM. EVF-SAM does not include elaborate designs or modules and is easy for scaling to larger models.
Training EVF-SAM is simple and conducted on referring segmentation datasets, e.g., RefCOCO [31], which is appropriate to adapt the original SAM for text prompts. Despite the simple architecture, our EVF-SAM achieves superior performance on referring expression segmentation tasks and outperforms previous attempts with Large Language Models [20, 21, 27], as shown in Fig. 1. The experimental results demonstrate that (1) using a multimodal encoder with the input text and image and (2) early fusion between the text and image contribute to the better-referring ability for SAM, showing a promising direction for text-prompted SAM. Additionally, the experiments also show the superiority of our EVF-SAM using a multimodal encoder over previous methods with decoder-only Large Language Models: (1) EVF-SAM reduces huge amounts of parameters, e.g., 82% parameters compared to LISA; (2) EVF-SAM relies less on handcrafted templates or instructions, which is more efficient and flexible; (3) EVF-SAM obtains better performance with less training data.
Figure 2: Comparisons of different Text-prompted SAM. (a) Given input texts, several works [28] leverage grounded detectors, e.g., Grounding DINO [32], to generate box prompts for SAM. (b) A natural idea to support text prompts is to use an off-the-shelf text encoder to generate text embeddings for SAM [1, 33] while the performance of referring segmentation is inferior. (c) Several works [20, 21, 27] adopt Large Language Models (LLM) or Large Multimodal Models (LMM) to generate prompt embeddings for SAM in an autoregressive manner, which incurs a large computation burden. (d) Our proposed EVF-SAM exploits an effective Multimodal Encoder for text-prompted SAM with higher performance and fewer parameters compared to LLM-based methods.
@@ 2.1 Text-Prompted Segment Anything Models
Segment Anything Model. SAM [1] is an interactive segmentation model capable of predicting non-semantic masks based on various types of prompts (points, boxes, coarse masks). Trained on a large-scale dataset, SAM demonstrates strong generalization capability for segmenting diverse common objects. Several works [2, 4] address the massive computation cost of SAM and propose efficient variants. Efficient-SAM [2] distils the image encoder of SAM, achieving comparable performance with significantly fewer parameters. Fast-SAM [4], leveraging the YOLOv8 [34] architecture, achieves a 50×50$\times$50 × speedup for inference. SAM-HQ[5] addresses the segmentation quality of SAM and utilizes low-level features from the image encoder to enhance the mask decoder for better accuracy. Although SAM excels in visual-based segmentation tasks with box/point/mask prompts, it currently lacks language understanding abilities and it’s infeasible to directly use text prompts for referring segmentation or semantic segmentation.
Text-Prompted explorations. Recently, several works [28, 4, 33] have explored text prompts for SAM to segment objects according to the instructions or referring expressions. Grounded-SAM [28] leverages the Grounding DINO [32] to obtain text-prompted boxes and feed the boxes to SAM for segmentation results, which formulates the non-end-to-end two-stage frameworks. Fast-SAM [4] matches the similarity of CLIP [29] features between the text and Region of Interest (RoI) of image. RefSAM [33] employs a lightweight cross-modal MLP to project the text embeddings of the referring expressions into SAM’s sparse embeddings and dense embeddings. LISA [20, 21] employs a Large Multimodal Model, e.g., LLaVA [35] to extract multimodal embeddings for SAM through the auto-regressive decoder. The aforementioned methods either suffer from poor performance or are computationally expensive. Referring expression segmentation based on SAM is a promising area for exploration, offering significant potential. We propose an effective end-to-end model that overcomes SAM’s limitations by enabling text-prompted segmentation capabilities.
@@ 2.2 Referring Expression Segmentation
Referring Segmentation via Text Encoders. Prevalent methods [13, 14, 15, 16] tend to leverage transformer-based text encoders, e.g., BERT [36] or CLIP [29], to encode expression texts into embeddings as guidance for segmentation. RefTr [37] uses a visual-language encoder to fuse image and text features and regresses the box and mask with a carefully designed query processor. LAVT [15] leverages a hierarchical Vision Transformer [38] (ViT) to perform language-aware visual encoding. CRIS [14] designs a vision-language decoder to merge CLIP features, propagating fine-grained semantic information from textual representations to each pixel-level activation. PolyFormer [16] follows the encoder-decoder structure, employing a transformer decoder to generate regression results. Novel methods pay attention to being compatible with multiple tasks to formulate a uniform model. UNINEXT [19], UniRef++ [17] and UniLSeg [18] employ similar frameworks but focus on utilizing datasets from different fields to empower their generalization capability. Although these traditional models are usually lightweight and achieve fine performance, They fail to integrate with large-scale foundation models, e.g., SAM[1], LLaVA[35], thereby struggling to keep pace with the trend of increasingly extensive pre-training.
TABLE I: Motivation analysis. Both CLIP and BEIT-3 are of Large scale, with comparable numbers of parameters. Specifically, CLIP has a total parameter count of 428M, while BEIT-3 totals 673M parameters. Metric of LLaVA [35] is borrowed from LISA-7B [20]
CLIP [29] (Text) CLIP [29] (Text+Image) BEIT-3 [36] (Text) BEIT-3 [30] (Text+Image) LLaVA [35] (Text+Image)
cIoU (RefCOCO) 63.4 67.9 65.1 83.7 79.1
Referring Segmentation via Large Language Models. In the context of the rapid development of Large Multimodal Models [35, 39, 40, 41] (LMM), a number of works [20, 21, 22, 23, 24, 25] have leveraged these models to encode expression texts for referring expression segmentation tasks. LISA [20, 21] finetune LLaVA [35] to make it able to answer questions related to segmentation with a fixed template like ”It is [SEG].”, where the hidden embeddings at the place of special token [SEG] will be seen as multimodal features extracted by LMM. PixelLM [22] extends LISA by building a segmentation codebook to enable multi-object segmentation. PerceptionGPT [23] proposes an end-to-end architecture. u-LLaVA [24] supports multi-task. PSALM [25] imports mask tokens to LMM input for better performance. However, those methods tend to adopt heavy architectures, especially the LLMs or LMMs, leading to a heavy computation burden for downstream applications. In contrast, we find that lightweight vision-language models perform better for encode text prompts for referring image segmentation.
@@ 3.1 Motivation: SAM with Vision-Language Models
Considering that SAM [1] has a strong generalization capability for image segmentation while the text-prompted ability has not been revealed, we investigate how to encode text prompts for SAM in this section. We started by using the vanilla text encoder, as shown in Fig. 3 and conducted preliminary experiments on RefCOCO (testA) to evaluate the referring ability of SAM, shown in Tab. I.
Figure 3: Architectural explorations for text-prompted SAM. ‘L’ and ‘V’ denote the text encoder and vision encoder. We mainly explore three schemes: (a) vanilla baseline with a simple text encoder, (b) multimodal inputs with a late fusion, i.e., concatenation, and (c) multimodal inputs with early vision-language fusions, e.g., BEIT-3 [30].
Multimodal referring information for SAM. SAM [1] has explored the feasibility of employing a CLIP text encoder to facilitate text-prompted segmentation, as illustrated in Fig. 3 (a). We owe its weak performance to the single-modal referring information. CLIP-prompted SAM achieves 63.4 cIoU at the RefCOCO/testA benchmark, far from well-defined baselines. CLIP exhibits strong alignment between text and image modalities, this alignment is insufficient for fine-grained tasks like segmentation. The referring information extractor should be provided with the input image and text prompts to ensure accurate alignment between the text expression and the relevant image region. We observe performance improvements after using multimodal prompts, i.e., 63.4 v.s. 67.9 for CLIP and 65.1 v.s. 83.7 for BEIT-3.

### v4

@@ Abstract
Segment Anything Model (SAM) has attracted widespread attention for its superior interactive segmentation capabilities with visual prompts while lacking further exploration of text prompts. In this paper, we empirically investigate what text prompt encoders (e.g., CLIP or LLM) are good for adapting SAM for referring expression segmentation and introduce the Early Vision-language Fusion-based SAM (EVF-SAM). EVF-SAM is a simple yet effective referring segmentation method which exploits multimodal prompts (i.e., image and text) and comprises a pre-trained vision-language model to generate referring prompts and a SAM for segmentation. Surprisingly, we observe that: (1) multimodal prompts and (2) vision-language models with early fusion (e.g., BEIT-3) are beneficial for prompting SAM for accurate referring segmentation. Our experiments show that the proposed EVF-SAM based on BEIT-3 can obtain state-of-the-art performance on RefCOCO/+/g for referring expression segmentation and demonstrate the superiority of prompting SAM with early vision-language fusion. In addition, the proposed EVF-SAM with 1.32B parameters achieves remarkably higher performance while reducing nearly 82% of parameters compared to previous SAM methods based on large multimodal models. Code and models will be made publicly available.
@@ 1 Introduction
Segment Anything Model (SAM) (Kirillov et al., 2023) brings interactive segmentation paradigm to public view. Well-trained on the SA-1B dataset, SAM achieves stunning performance and quickly becomes popular as a vision foundation model for object localization and beyond. Various SAM variants (Xiong et al., 2023; Zhang et al., 2023; Zhao et al., 2023; Ke et al., 2024) have been explored, achieving better efficiency or higher precision. Despite SAM’s surprising abilities like point-prompted and box-prompted segmentation, it is a pity that the text-prompted segmentation ability remains conceptual. We retrospect such task to Referring Expression Segmentation (RES). RES focuses on the solution that one predicts the segmentation mask according to the text description given by users, which enjoys several explorations by some traditional end-to-end models (Hu et al., 2016; Liu et al., 2017; Shi et al., 2018; Chen et al., 2019; Ye et al., 2019; Hu et al., 2020; Ding et al., 2021; Li & Sigal, 2021b; Wang et al., 2022b; Yang et al., 2022; Liu et al., 2023c; Wu et al., 2023; Liu et al., 2023e; Yan et al., 2023), and is broadened by some Large Multimodal Models (LMM) (Lai et al., 2023; Yang et al., 2023; Ren et al., 2023; Pi et al., 2023; Xu et al., 2023; Zhang et al., 2024; Xia et al., 2023; Rasheed et al., 2023).  [^1] Equal contribution.  [^2] Corresponding author (xgwang@hust.edu.cn).
Figure 2: Comparisons of different Text-prompted SAM. (a) Given input texts, several works (Ren et al., 2024) leverage grounded detectors, e.g., Grounding DINO (Liu et al., 2023d), to generate box prompts for SAM. (b) A natural idea to support text prompts is to use an off-the-shelf text encoder to generate text embeddings for SAM (Kirillov et al., 2023; Li et al., 2023) while the performance of referring segmentation is inferior. (c) Several works (Lai et al., 2023; Yang et al., 2023; Rasheed et al., 2023) adopt Large Language Models (LLM) or Large Multimodal Models (LMM) to generate prompt embeddings for SAM in an autoregressive manner, which incurs a large computation burden. (d) Our proposed EVF-SAM exploits an effective Multimodal Encoder for text-prompted SAM with higher performance and fewer parameters compared to LLM-based methods.
The key challenge lies in empowering SAM with language understanding ability for segmentation according to text prompts, e.g., referring expression segmentation. Fig. 2 summarizes previous works which explore the text-prompted abilities of SAM: (a) SAM with grounded detector: A two-stage framework where a grounded detector generates a bounding box to prompt SAM, e.g., Grounded-SAM (Ren et al., 2024). However, those methods suffer from a sub-optimal architecture, where segmentation heavily relies on the accuracy of the detector, and it is difficult to optimize due to its non-end-to-end nature. (b) SAM with text encoder: A off-the-shelf text encoder, e.g., CLIP (Radford et al., 2021), is used to encode the text prompt, providing text embeddings for SAM. Whereas the semantic gap exists between the text embeddings and SAM which is pre-trained with geometric prompts, i.e., points or boxes, thus the segmentation performance is inferior. (c) SAM with LLM: A Large Language Model (LLM) (or Large Multimodal Model) is employed and fine-tuned to get the desired embeddings about object information. The embeddings will be used to predict segmentation masks based on image features. However, these LLM-based models are often computationally expensive, requiring massive memory and computation budgets, and the training is challenging. Additionally, complex conversation templates need to be manually designed to instruct the LLM for referring segmentation. Can we leverage a more efficient but effective method to empower SAM with text-prompted ability in an end-to-end manner?
@@
Motivated by the above observations, we extend SAM for language understanding and text-prompt capabilities by incorporating a Multimodal Encoder with Early Vision-Language Fusion (EVF) and present EVF-SAM in this paper. The proposed EVF-SAM aims to be a simple framework to prompt SAM with texts and illustrate how to prompt SAM to follow referring expressions effectively. EVF-SAM is built on the off-the-shelf foundation models and comprises a Multimodal Encoder, an early-fused vision-language model, e.g., BEIT-3 (Wang et al., 2022a), and a simple projector to generate prompt embeddings for SAM. EVF-SAM does not include elaborate designs or modules and is easy for scaling to larger models.
Training EVF-SAM is simple and conducted on referring segmentation datasets, e.g., RefCOCO (Yu et al., 2016), which is appropriate to adapt the original SAM for text prompts. Despite the simple architecture, our EVF-SAM achieves superior performance on referring expression segmentation tasks and outperforms previous attempts with Large Language Models (Lai et al., 2023; Yang et al., 2023; Rasheed et al., 2023), as shown in Fig. 1. The experimental results demonstrate that (1) using a multimodal encoder with the input text and image and (2) early fusion between the text and image contribute to the better-referring ability for SAM, showing a promising direction for text-prompted SAM. Additionally, the experiments also show the superiority of our EVF-SAM using a multimodal encoder over previous methods with decoder-only Large Language Models: (1) EVF-SAM reduces huge amounts of parameters, e.g., 82% parameters compared to LISA; (2) EVF-SAM relies less on handcrafted templates or instructions, which is more efficient and flexible; (3) EVF-SAM obtains better performance with less training data.
@@ 2.1 Text-Prompted Segment Anything Models
Segment Anything Model. SAM (Kirillov et al., 2023) is an interactive segmentation model capable of predicting non-semantic masks based on various types of prompts (points, boxes, coarse masks). Trained on a large-scale dataset, SAM demonstrates strong generalization capability for segmenting diverse common objects. Several works (Xiong et al., 2023; Zhao et al., 2023) address the massive computation cost of SAM and propose efficient variants. Efficient-SAM (Xiong et al., 2023) distils the image encoder of SAM, achieving comparable performance with significantly fewer parameters. Fast-SAM (Zhao et al., 2023), leveraging the YOLOv8 (Jocher et al., 2023) architecture, achieves a 50×50$\times$50 × speedup for inference. SAM-HQ (Ke et al., 2024) addresses the segmentation quality of SAM and utilizes low-level features from the image encoder to enhance the mask decoder for better accuracy. Although SAM excels in visual-based segmentation tasks with box/point/mask prompts, it currently lacks language understanding abilities and it’s infeasible to directly use text prompts for referring segmentation or semantic segmentation.
Text-Prompted explorations. Recently, several works (Ren et al., 2024; Zhao et al., 2023; Li et al., 2023) have explored text prompts for SAM to segment objects according to the instructions or referring expressions. Grounded-SAM (Ren et al., 2024) leverages the Grounding DINO (Liu et al., 2023d) to obtain text-prompted boxes and feed the boxes to SAM for segmentation results, which formulates the non-end-to-end two-stage frameworks. Fast-SAM (Zhao et al., 2023) matches the similarity of CLIP (Radford et al., 2021) features between the text and Region of Interest (RoI) of image. RefSAM (Li et al., 2023) employs a lightweight cross-modal MLP to project the text embeddings of the referring expressions into SAM’s sparse embeddings and dense embeddings. LISA (Lai et al., 2023; Yang et al., 2023) employs a Large Multimodal Model, e.g., LLaVA (Liu et al., 2023b) to extract multimodal embeddings for SAM through the auto-regressive decoder. The aforementioned methods either suffer from poor performance or are computationally expensive. Referring expression segmentation based on SAM is a promising area for exploration, offering significant potential. We propose an effective end-to-end model that overcomes SAM’s limitations by enabling text-prompted segmentation capabilities.
@@ 2.2 Referring Expression Segmentation
Referring Segmentation via Text Encoders. Prevalent methods (Li & Sigal, 2021b; Wang et al., 2022b; Yang et al., 2022; Liu et al., 2023c) tend to leverage transformer-based text encoders, e.g., BERT (Devlin et al., 2018) or CLIP (Radford et al., 2021), to encode expression texts into embeddings as guidance for segmentation. RefTr (Li & Sigal, 2021a) uses a visual-language encoder to fuse image and text features and regresses the box and mask with a carefully designed query processor. LAVT (Yang et al., 2022) leverages a hierarchical Vision Transformer (Dosovitskiy et al., 2020) (ViT) to perform language-aware visual encoding. CRIS (Wang et al., 2022b) designs a vision-language decoder to merge CLIP features, propagating fine-grained semantic information from textual representations to each pixel-level activation. PolyFormer (Liu et al., 2023c) follows the encoder-decoder structure, employing a transformer decoder to generate regression results. Novel methods pay attention to being compatible with multiple tasks to formulate a uniform model. UNINEXT (Yan et al., 2023), UniRef++ (Wu et al., 2023) and UniLSeg (Liu et al., 2023e) employ similar frameworks but focus on utilizing datasets from different fields to empower their generalization capability. Although these traditional models are usually lightweight and achieve fine performance, They fail to integrate with large-scale foundation models, e.g., SAM(Kirillov et al., 2023), LLaVA(Liu et al., 2023b), thereby struggling to keep pace with the trend of increasingly extensive pre-training.
Referring Segmentation via Large Language Models. In the context of the rapid development of Large Multimodal Models (Liu et al., 2023b; Bai et al., 2023; Sun et al., 2023b; a) (LMM), a number of works (Lai et al., 2023; Yang et al., 2023; Ren et al., 2023; Pi et al., 2023; Xu et al., 2023; Zhang et al., 2024) have leveraged these models to encode expression texts for referring expression segmentation tasks. LISA (Lai et al., 2023; Yang et al., 2023) finetune LLaVA (Liu et al., 2023b) to make it able to answer questions related to segmentation with a fixed template like ‘It is [SEG].’, where the hidden embeddings at the place of special token [SEG] will be seen as multimodal features extracted by LMM. PixelLM (Ren et al., 2023) extends LISA by building a segmentation codebook to enable multi-object segmentation. PixelLLM (Xu et al., 2024) empowers the vision-language model to take locations (e.g., a set of points or boxes) as either inputs or outputs. PerceptionGPT (Pi et al., 2023) proposes an end-to-end architecture. u-LLaVA (Xu et al., 2023) supports multi-task. PSALM (Zhang et al., 2024) imports mask tokens to LMM input for better performance. However, those methods tend to adopt heavy architectures, especially the LLMs or LMMs, leading to a heavy computation burden for downstream applications. In contrast, we find that lightweight vision-language models perform better for encode text prompts for referring image segmentation.
@@ 3.1 Motivation: SAM with Vision-Language Models
Figure 3: Architectural explorations for text-prompted SAM. ‘L’ and ‘V’ denote the text encoder and vision encoder. We mainly explore three schemes: (a) vanilla baseline with a simple text encoder, (b) multimodal inputs with a late fusion, i.e., concatenation, and (c) multimodal inputs with early vision-language fusions, e.g., BEIT-3 (Wang et al., 2022a).
Considering that SAM (Kirillov et al., 2023) has a strong generalization capability for image segmentation while the text-prompted ability has not been revealed, we investigate how to encode text prompts for SAM in this section. We started by using the vanilla text encoder, as shown in Fig. 3 and conducted preliminary experiments on RefCOCO (testA) to evaluate the referring ability of SAM, shown in Tab. 1.
Multimodal referring information for SAM. SAM (Kirillov et al., 2023) has explored the feasibility of employing a CLIP text encoder to facilitate text-prompted segmentation, as illustrated in Fig. 3 (a). We owe its weak performance to the single-modal referring information. CLIP-prompted SAM achieves 63.4 cIoU at the RefCOCO/testA benchmark, far from well-defined baselines. CLIP exhibits strong alignment between text and image modalities, this alignment is insufficient for fine-grained tasks like segmentation. The referring information extractor should be provided with the input image and text prompts to ensure accurate alignment between the text expression and the relevant image region. We observe performance improvements after using multimodal prompts, i.e., 63.4 v.s. 67.9 for CLIP and 65.1 v.s. 83.7 for BEIT-3.
Table 1: Motivation analysis. Both CLIP and BEIT-3 are of Large scale, with comparable numbers of parameters. Specifically, CLIP has a total parameter count of 428M, while BEIT-3 totals 673M parameters. Metric of LLaVA (Liu et al., 2023b) is borrowed from LISA-7B (Lai et al., 2023)
CLIP (Text) CLIP (Text+Image) BEIT-3 (Text) BEIT-3 (Text+Image) LLaVA (Text+Image)
cIoU (RefCOCO) 63.4 67.9 65.1 83.7 79.1
Early-fused architecture. Some existing works, e.g., LISA (Lai et al., 2023; Yang et al., 2023), UniLSeg (Liu et al., 2023e), advocate for fusing visual and textual information simply before the mask generator and are widely considered as ‘early fusion’. However, we argue that these approaches are not early enough. As illustrated in Fig. 3 (b), we define such fusion for separately encoded single-modal prompts as ‘late fusion’. In contrast, as shown in Fig. 3 (c), we define the fusion during feature extraction, where both modalities can access the dense information of the other one, as ‘early fusion’, e.g., ViLT (Kim et al., 2021), BEIT-3 (Wang et al., 2022a), which incorporate the cross-modal fusions within the encoder.
We leverage the ‘early-fusion’ vision-language model as the Multimodal Encoder to generate prompt embeddings for SAM. Tab. 1 shows that our investigation indicates that early-fusion outperforms late-fusion, i.e., 83.7 for BEIT-3 and 67.9 for CLIP. We believe the early-fused architecture, as defined by our approach, is beneficial for encoding text prompts since the cross-modal fusions will further enhance the semantic representation for text embeddings. In addition, the text-to-image fusions guide the image branch to aggregate features which are aligned with text prompts, making the output embeddings more accurate for prompt SAM.
Encoder-based feature extractor. Recently, LISA (Lai et al., 2023; Yang et al., 2023) and several LLM-based methods (Rasheed et al., 2023; Ren et al., 2023) acquire the prompt embeddings for SAM with a special token through the auto-regressive generation. However, the uncontrolled length of the answering query introduces instability during both training and inference. Forcing the model to conform to a specific answering template can lead to language drift. In contrast, encoder-based architectures can maintain a consistent sequence length of inputs and outputs. Utilizing the encoder-based method not only offers convenience but also yields superior performance, i.e., 79.1 for LLaVA and 83.7 for BEIT-3. Notably, the encoder-based text-prompted SAM will reduce a massive computation burden compared to the LLM-based methods.
3.2 Architecture